Circular No. 7956

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 Q1 (VAN NESS) Michael E. Van Ness, Lowell Observatory, reports his discovery of a

comet with a moderately condensed 24′′ coma and a fan-shaped tail ∼ 200′′ long in p.a. 270^{o} on LONEOS images taken on Aug. 17. The cometary nature was confirmed by several other CCD astrometric observers on Aug. 18, who all reported a generally diffuse coma with diameter ∼ 14′′–25′′; several also noted a tail ∼ 25′′–120′′ long in p.a. 260^{o}–270^{o}.

2002 UT  α_{2000}  δ_{2000}  m_{1}
Aug. 17.44384  3^{h}57^{m}42.^{s}77  +1^{o}08′26′′.4  16.9

The available astrometric data (including observer information) and the following preliminary orbital elements appear on MPEC 2002-Q08.

T = 2002 July 25.502 TT  ω = 190.^{o}673
Ω = 178.447 }2000.0
q = 1.65463 AU  i = 38.671

SUPERNOVA 2002en IN UGC 12289 Further to IAUC 7906 and 7953, W. D. Li reports the LOTOSS dis-

covery of an apparent supernova (mag ∼ 18.3) on unfiltered CCD images taken with KAIT on Aug. 18.3 and 19.3 UT. SN 2002en is located at α = 22^{h}59^{m}40^{s}.82, δ = +24^{o}04′17′′.5 (equinox 2000.0), which is 9′′.9 west and 12′′.2 south of the nucleus of UGC 12289. A KAIT image taken on Aug. 12.3 showed nothing at this position (limiting mag ∼ 19.0).

COMETS C/2002 N2, C/2002 P2, AND C/2002 P3 (SOHO) Further to IAUC 7952, D. Hammer reports measurements for three

comets found on SOHO website images by R. Kracht. C/2002 P2 (a Kreutz comet) was seen on C3 coronagraph images; the others were visible only on C2 images; C/2002 P3 is a member of the Meyer group (cf. IAUC 7832), and C/2002 N2 of the Kracht group (IAUC 7952). Reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  2002 UT  α_{2000}  δ_{2000}  MPEC
C/2002 N2 July 11.646  7^{h}15.^{m}8  +22^{o}31′  2002-Q02
C/2002 P2 Aug. 10.513  8 59.6  +13 58  2002-P55
C/2002 P3  12.504  9 21.7  +16 19  2002-P63
2002 August 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green